OpenStudy (anonymous):

if I have an expression (x+(x^3/3)) how can I factor out the 1/3 and keep the X in parenthesis

OpenStudy (anonymous):

\[x+\frac{x^3}{3}\] this?

OpenStudy (anonymous):

assuming that was your problem \[\frac{3x}{3}+\frac{x^3}{3}\] can you factor out a 1/3 now?